two sphere just fit into a cylindrical container of radius 4.5 cm .
Calculate the volume occupied by both sphere.

Answer by mananth(16949)

Volume of sphere = V= 4/3 * pi* r^3
V= 4/3 * 3.14*(4.5)^3
V= 381.5 cu.cm of i sphere
2*381.5 = 763 cu.cm Volume of two spheres